David Oliver and Umar Ansari, with capital balances of $58,000 and $78,000, respectively, decide to liquidate their partnership. After selling the noncash assets and paying the liabilities, there is $170,000 of cash remaining. If the partners share income and losses equally, how should the cash be distributed?
Answer #1

Capital, David = $58,000

Capital, Umar = $78,000

Cash remaining after paying liabilities = $170,000

Gain on disposal of assets = Cash remaining after paying liabilities - Capital, Umar - Capital, David

= 170,000 - 78,000 - 58,000

= $34,000

Gain to be shared by David = 34,000 x 1/2

= $17,000

Gain to be shared by Umar = 34,000 x 1/2

= $17,000

Cash received by David as final distribution from liquidation of partnership = Capital, David + Share of gain

= 58,000 + 17,000

= $75,000

Cash received by Umar as final distribution from liquidation of partnership = Capital, Umar + Share of gain

= 78,000 + 17,000

= $95,000

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.